Use Intermediate Value Theorem lial p(x) = x4 +7x = 9 has two real root. 8. df Open with Google Docs Then use your calculatorPlease do 10 & 11
Answer #1

Than 0 > Suppese <S e)-prx=JT- SL) Sunce 20 ,20 ) hen yentePle al nawmber o Such te xelo, fDO) retipi furdion Heme pP=1YoX h-b-a h= L Myp n sbityal 44+16+16- 36 16 16 9424t16 49 1+200+1 =다 PCi) Supp nr- |월 후구W ddu-u T5 mgmpP m. Sylength of yal . i+ mdt -P,)= m +25+36 16 126 NR IF Pird eryals -UP) fz+FG&3 + = -+T%S